OpenStudy (anonymous):

The speed of a cycle increases from 2.5 m/s to 5 m/s. The total mass of the cyclist including the cycle is 120 kg. Calculate the work done by the cyclist to increase the speed.

OpenStudy (shamim):

i m trying to solve this problem

OpenStudy (shamim):

here u=2.5 v=5 m=120kg work done W=?

OpenStudy (shamim):

\[W=FS=maS\]

OpenStudy (shamim):

again we know\[v ^{2}=u ^{2}+2aS\]

OpenStudy (shamim):

\[or, 2aS=v ^{2}-u ^{2}\]

OpenStudy (shamim):

\[or, aS=\frac{ v ^{2}-u ^{2} }{ 2 }\]

OpenStudy (shamim):

putting this value of aS in the 1st equation we get

OpenStudy (shamim):

\[W=m(\frac{ v ^{2}-u ^{2} }{ 2 })\]

OpenStudy (shamim):

\[or, W=120 \times (\frac{ 5^{2}-2.5^{2} }{ 2 })\]

OpenStudy (shamim):

or, W=1125 joule

OpenStudy (shamim):

r u following me

OpenStudy (vincent-lyon.fr):

Use work-energy theorem, and write: \(W=m(\Large\frac{ v ^{2}-u ^{2} }{ 2 })\) directly. This is more general than the derivation provided by shamim, as it will be true even for non-uniform acceleration. But this will only be true if the cyclist is on a horizontal road.